Consider the inventory, sales, and COGS of Costco and Walmart:

(units are $ Millions)

Costco

Walmart

Inventories

3,643

28,447

Sales (net)

48,106

286,103

COGS

41,651

215,493

a) How many days does a product stay in Costco’s inventory before it is sold? Assume stores operate 365 days per year.

b) Suppose Costco and Walmart carry the same household cleaner. How much lower is the inventory cost for this item at Costco than at Walmart (in terms of percentages) assuming that the companies source this item from the same wholesaler?
Answer #1

(a) Days in inventory = Inventory cost / COGS = 3643/41651

Days in Inventory (Costco) = 0.0875 years = 31.92 days

(b) Inventory Cost =item cost*Days in inventory*holding cost rate

because item cost and holding cost rate are same, therefore, inventory cost is directly dependent on days of inventory

At Costco --- 0.0875

Days in Inventory (Walmart) = 28447/ 215493 = 0.132 years

At Walmart --- 0.132

Difference in %= (0.1320-0.0875)/0.1320 * 100 = 33.71%